Please help me solve this equation:
{x^2+x+sqrt(x^2+x)-2=0}

Answer by mananth(16946)

{x^2+x+sqrt(x^2+x)-2=0}
let sqrt(x^2+x) be a
a^2+a-2)=0
a^2+2a-a-2=0
a(a+2)-1(a+2)=0
(a+2)(a-1)=0
(x^2+x+2)(x^2+x-1)=0
Find x1,x2 by quadratic formula method for both